# BLIPs: Bayesian Learned Interatomic Potentials

### v1

@@ A.1 KL-divergence proof
In practice, we weight the KL term by a factor $\lambda$ as normally done for ELBO optimisation (Alemi et al.,, 2018; Higgins et al.,, 2017).
@@ A.2 Predictive Distribution and Uncertainty Estimates
In practice, this corresponds to performing multiple stochastic forward passes through the network and averaging the outputs, following the standard Monte Carlo approximation of predictive distributions (Gal and Ghahramani,, 2016). Alternatively, a deterministic approximation can be used by evaluating the network at the posterior mean weights when uncertainty is not needed.
@@ B.1 NBody System
We utilise the dataset introduced by Fuchs et al., (2020), which consists of simulations of five particles interacting in a three-dimensional box under Coulomb’s law. This dataset is publicly available at: https://github.com/FabianFuchsML/se3-transformer-public/tree/master/experiments/nbody/data_generation. Following the setup of Satorras et al., (2021), we generate the dataset using the parameters --num-train 10000 --seed 43 --sufix small. We follow the protocol of Satorras et al., (2021) and use $3000$ samples each for training, validation, and testing, ensuring that there is no overlap between the splits.
@@ B.2 Ammonia and Silica Glass Systems
For the ammonia system, we utilize the dataset introduced by Tan et al., (2023), which comprises a small number of NH3 molecular geometries sampled across a range of energy configurations. The dataset is publicly available at: https://archive.materialscloud.org/records/tz6a0-hsb25. We follow the same training/validation/test split protocol as Tan et al., (2023).
For the silica glass system, we also use the dataset from Tan et al., (2023), which consists of large SiO2 structures extracted from diverse, long molecular dynamics trajectories. These trajectories span a range of physical conditions, including temperature equilibration, cooling, uniaxial tension, compression, and more. The dataset is publicly available at: https://archive.materialscloud.org/records/tz6a0-hsb25. We adopt the same validation/test split protocol as Tan et al., (2023), and vary the size of the training set ($32$, $128$, and $1024$ samples).
Additional details regarding the DFT calculations used to generate both datasets are provided in the Supplementary Information of Tan et al., (2023).
@@
We follow the same experimental procedure as Tan et al., (2023): for prediction we compute the mean absolute error between predicted energy and forces $\hat{E},\bm{\hat{F}}$ against true ones:
@@
To evaluate uncertainty, we compute the Expected Calibration Error (ECE) on forces, following the approach proposed by Tan et al., (2023). This choice is motivated by the fact that forces are more sensitive to overfitting and distributional shift, and thus provide a more informative measure of epistemic uncertainty than energy-based metrics. We estimate uncertainties using the mean force per structure, denoted by $\bar{F}_{i}$ for structure $i$, which is obtained by averaging the per-atom forces and flattening across the spatial dimensions. The corresponding standard deviation, $\bar{\sigma}_{i}$, is computed as the square root of the average variance norm per atom, following the procedure in Tan et al., (2023). The ECE is then computed as:
@@
where $\mathbb{Q}$ is the gaussian cumulative density function. The integral is computed by trapezoidal rule. We further evaluate the rank correlation between predicted variance $\bar{\sigma}_{i}^{2}$ and squared error norm per atom using the Spearman correlation coefficients using the Scipy (Virtanen et al.,, 2020) implementation.
@@ Appendix C Architectures and Training Details
All our models have been trained on a single 64GB A-100 Nvidia GPU. We report in Table 4 the hyperparameters used across experiments shared among different models.
Table 4: Summary of main training-related hyperparameters for different experiments. These hyperparameters are shared among different models. Adam is the optimiser in Kingma and Ba, (2014), while AdamW the one in Loshchilov and Hutter, (2017).
Hyper-parameter NBody Ammonia Silica
Optimization Epochs 10000 500 50
Batch Size 100 64 8
Optimiser Adam Adam AdamW
Loss MSE MAE HuberLoss(delta=0.01)
Maximum learning rate 5e45e-4 1e31e-3 1e51e-5
Learning rate scheduling Constant ReduceLROnPlateau Cosine with Warmup
Gradient Clip Norm Value NA NA 0.5
Warmup Iteration Phase NA NA 5%
@@ C.1 NBody System
For the EGNN baseline, we employ the velocity variant from Satorras et al., (2021), where the initial velocity norm of each particle is mapped through a linear layer and included in the initial node features $\mathbf{h}^{0}_{i}$. Particle charges are again encoded as edge attributes via the pairwise product $a_{ij}=c_{i}c_{j}$. The message and update functions have the same parameterisation as in the GNN model. Following Satorras et al., (2021), no regression head is used, as the network directly predicts particle positions.
Regarding the baseline methods, for MC Dropout we follow Gal and Ghahramani, (2016) and apply dropout with varying probabilities after all linear layers of the message-passing network. For Deep Ensemble, we train the same architecture four times with different random initialisation seeds.
Finally, for BLIP, we employ a compact invariant posterior network consisting of two linear layers of size $64$ with swish activations, used for both the message and update coefficients. Dropout probabilities $p$ are predicted and converted to coefficients via p/(1p)p/(1-p). The message inference network head takes as input the edge attribute $a_{ij}$ together with $\mathbf{h}^{0}_{i}$ and $\mathbf{h}^{0}_{j}$, whereas the update inference network head only takes $\mathbf{h}^{0}$. For stability, we clamp the maximum dropout probability to $0.8$. During training, we use a prior dropout probability of $0.5$ and weight the regularisation term $D_{\mathrm{KL}}$ by $\lambda=0.01$.
@@ C.2 Ammonia Molecule System
We adopt the PaiNN architecture (Schütt et al.,, 2021) as our machine learning interatomic potential. To ensure a fair comparison, we use the same hyperparameter settings as in Tan et al., (2023), which are publicly available at: https://github.com/learningmatter-mit/UQ_singleNN.
@@
The message inference network head takes as input the squared interatomic distance $\|\bm{r}_{i}-\bm{r}_{j}\|^{2}$, whereas the update inference network head takes only an embedding of the element atomic number $Z$. Additional architecture and training details are provided in Section C. For stability, we clamp the maximum dropout probability to $0.8$. During training, we use a prior dropout probability of $0.5$ and weight the regularisation term $D_{\mathrm{KL}}$ by $\lambda=10$.
@@ C.3 Silica Glass System
We adopt the ORB v3 architecture (Rhodes et al.,, 2025) as our machine learning interatomic potential, pretrained on ab initio molecular dynamics data from OMAT 24 (Barroso-Luque et al.,, 2024). The Deep Ensemble variant finetunes four different models with different initialisation seeds.
@@
The edge network takes as input a concatenation of the pairwise distance vector (normalised to unit length) and 20 Gaussian radial basis functions applied to the raw edge distance. The node network processes atomic number embeddings. Finally, dropout probabilities $p$ are predicted and transformed into multiplicative noise coefficients using the expression p/(1p)p/(1-p). For stability, we clamp $p$ to the range $[0.00001,0.1]$. During training, we use a prior dropout probability of $0.01$ and weight the regularisation term $D_{\mathrm{KL}}$ by $\lambda=0.001$. The lower clamp values, $\lambda$, and reduced prior dropout are chosen because we find that the original ORB model weights are highly sensitive to even minimal perturbations, likely because they were not trained with dropout or other weight-perturbation techniques. In addition, we use a different learning rate (1e41e-4) for the inference network to enable faster adaptation compared to the pretrained weights.
@@ Appendix D Algorithms
Local-reparametrization trick:
The local-reparametrization trick (Kingma et al.,, 2015) was introduced to efficiently sample the predictive distribution of Bayesian linear layers. For a single linear layer with input $\bm{h}$ and weight matrix $\bm{\omega}\sim\mathcal{N}(\bm{\theta},\alpha\bm{\theta}^{2})$, the output $\bm{z}$ also follows a normal distribution:
$\bm{z}\sim\mathcal{N}(\bm{\gamma},\mathrm{diag}(\bm{\delta})),$ (21)
where each output element $z_{j}$ can be written as
with
This shows that instead of sampling weights $\bm{\omega}$ directly, one can sample the output $\bm{z}$ by sampling from a Gaussian with analytically computed mean and variance, which reduces variance and computational cost. Below the algorithmic version.
[…]
4:return $\bm{m}+\bm{std}\odot\bm{eps}$
In standard deep learning libraries, the above algorithm can be easily implemented by overriding the linear layer module. In our source code, we provide built-in wrappers that implement this functionality for convenience and ease of use.
@@
12:  $\mathcal{L}\leftarrow\lambda_{\rm{E}}\text{MAE}_{\rm{E}}+\lambda_{\rm{F}}\text{MAE}_{\rm{F}}+\lambda D_{\mathrm{KL}}(\bm{\alpha},\bm{\beta})$ {Loss includes MAEs and KL divergence term from Eq. (10)}
@@
Uncertainty quantification is performed via multiple stochastic forward passes through the model using Monte Carlo (MC) sampling. Importantly, the inference network is evaluated only once to extract the variational parameters $\bm{\alpha},\bm{\beta}$, which are reused across MC samples. This amortization significantly reduces computational overhead. While more accurate moment-matching approximations could be used for the output distributions, we leave this for future work.
@@
If only predictive means are needed and uncertainty estimates are not required, a single deterministic forward pass can be performed using the mean of the variational distributions (i.e., without stochastic sampling), effectively matching the performance speed of a standard MLIP.

### v2

@@ A.1 KL-divergence proof
In practice, we weight the KL term by a factor $\lambda$ as normally done for ELBO optimisation (Alemi et al., 2018; Higgins et al., 2017).
@@ A.2 Predictive Distribution and Uncertainty Estimates
In practice, this corresponds to performing multiple stochastic forward passes through the network and averaging the outputs, following the standard Monte Carlo approximation of predictive distributions (Gal and Ghahramani, 2016). Alternatively, a deterministic approximation can be used by evaluating the network at the posterior mean weights when uncertainty is not needed.
@@
A.3 BLIP Stochastic Equivariance and Local Reparametrization
This section is dedicated to the formal proofs of the equivariance in the BLIP model. We first provide a general proposition for Stochastic Equivariant Maps, and we later connect this to the BLIP update rule.
Stochastic Equivariant Maps:
Let $G$ be a group acting measurably on spaces $X$ and $Y$. Let $g\cdot x\in X$, and $g\cdot y\in Y$ denote the respective actions. Assume:
  1. 1.
    An invariant map $\mathrm{inv}:X\to\mathrm{Inv}(X)$ such that
    $\mathrm{inv}(g\cdot x)=\mathrm{inv}(x)\quad\forall g\in G,\,x\in X.$
  2. 2.
    A class $\mathcal{F}$ of deterministic equivariant maps $f:X\to Y$, i.e.
    $f(g\cdot x)=g\cdot f(x)\quad\forall f\in\mathcal{F},\,g\in G,\,x\in X.$
  3. 3.
    A measurable map
    $E:\mathrm{Inv}(X)\to\mathcal{P}(\mathcal{F})$
    assigning to each invariant value a probability distribution over equivariant maps.
Define a stochastic map:
$P[y\mid x]=\int_{\mathcal{F}}\delta(y-f(x))dE(\mathrm{inv}(x))$ (15)
i.e. sample $f\sim E(\mathrm{inv}(x))$ and output the deterministic value $f(x)$.
Proposition 1.
The stochastic map $S$ is $G$-equivariant:
$P[y\mid g\cdot x]=g_{*}P[y\mid x]\quad\forall g\in G,\,x\in X,$ (16)
where $g_{*}:\mathcal{P}(Y)\to\mathcal{P}(Y)$ is the pushforward by the action $y\mapsto g\cdot y$.
Proof.
Fix $g\in G$ and $x\in X$. By definition of $S$,
$P[y\mid g\cdot x]=\int_{\mathcal{F}}\delta\!\bigl(y-f(g\cdot x)\bigr)\,dE(\mathrm{inv}(g\cdot x)).$ (17)
Since $\mathrm{inv}$ is invariant, $\mathrm{inv}(g\cdot x)=\mathrm{inv}(x)$. Thus
$P[y\mid g\cdot x]=\int_{\mathcal{F}}\delta\!\bigl(y-f(g\cdot x)\bigr)\,dE(\mathrm{inv}(x)).$ (18)
Because each $f\in\mathcal{F}$ is $G$-equivariant, $f(g\cdot x)=g\cdot f(x)$. Substituting,
$P[y\mid g\cdot x]=\int_{\mathcal{F}}\delta\!\bigl(y-g\cdot f(x)\bigr)\,dE(\mathrm{inv}(x)).$ (19)
Therefore, by the definition of pushforward $g_{*}\delta(\,\cdot-f(x))\;=\;\delta(\,\cdot-g\cdot f(x))$ if follows:
$P[y\mid g\cdot x]=g_{*}\left(\int_{\mathcal{F}}\delta\!\bigl(y-f(x)\bigr)\,dE(\mathrm{inv}(x))\right)=g_{*}P[y\mid x].$ (20)
Corollary 1.
The BLIP update in equation 7 rule applied to an equivariant MLIP yields equivariant stochastic MLIPs.
To clearly see why Corollary 1 holds, we highlight the following equivalence between the BLIP update and Proposition 1:
  • The invariant map $\mathrm{inv}$, is equivalent to the BLIP inference network $E$, mapping the input to the invariant variational dropout coefficients (equation 6).
  • The equivariant map $f$ is an equivariant MLIP parametrized by parameters $\theta$.
  • The stochastic map $P[y\mid x]$ is the once induced by the BLIP update (equation 7).
It is important to notice that the above corollary applies to any family of equivariant neural networks, including UMA (Wood et al., 2025) and MACE (Batatia et al., 2022). In what follows, we show that the local reparameterization trick (sampling activations) breaks equivariance, whereas perturbing the weights, in principle does not, opening further research to sample activations while remaining equivariant efficiently.
Local-reparametrization trick:
[…]
4:return $\bm{m}+\bm{std}\odot\bm{eps}$
The local-reparametrization trick (Kingma et al., 2015) was introduced to sample the predictive distribution of Bayesian linear layers efficiently. For a single linear layer with input $\bm{h}$ and weight matrix $\bm{\omega}\sim\mathcal{N}(\bm{\theta},\alpha\bm{\theta}^{2})$, the output $\bm{z}$ also follows a normal distribution:
$\bm{z}\sim\mathcal{N}(\bm{\gamma},\mathrm{diag}(\bm{\delta})),$ (21)
where each output element $z_{j}$ can be written as
with
This shows that instead of sampling weights $\bm{\omega}$ directly, one can sample the output $\bm{z}$ by sampling from a Gaussian with analytically computed mean and variance, which reduces variance and computational cost. In standard deep learning libraries, the above algorithm can be easily implemented by overriding the linear layer module (see Algorithm 1).
Local-reparametrization trick for equivariant linear maps:
When applying the local reparametrization trick, we push the noise from the weights to the activation, reducing the computational cost. If we apply this to a vector of irreducible dimensions, the trick introduces a different noise scale for each coordinate of the irrep. Since an irrep transforms by mixing its coordinates through a fixed linear representation, all components must be scaled uniformly to preserve equivariance. Local reparam instead applies independent variances to each channel, breaking the required coupling across the irrep. Thus, the resulting stochastic activation is no longer equivariant, despite the underlying weight distribution being equivariant.
@@ B.1 NBody System
We utilise the dataset introduced by Fuchs et al. (2020), which consists of simulations of five particles interacting in a three-dimensional box under Coulomb’s law. This dataset is publicly available at: https://github.com/FabianFuchsML/se3-transformer-public/tree/master/experiments/nbody/data_generation. Following the setup of Satorras et al. (2021), we generate the dataset using the parameters --num-train 10000 --seed 43 --sufix small. We follow the protocol of Satorras et al. (2021) and use $3000$ samples each for training, validation, and testing, ensuring that there is no overlap between the splits.
@@ B.2 Ammonia and Silica Glass Systems
For the ammonia system, we utilize the dataset introduced by Tan et al. (2023), which comprises a small number of NH3 molecular geometries sampled across a range of energy configurations. The dataset is publicly available at: https://archive.materialscloud.org/records/tz6a0-hsb25. We follow the same training/validation/test split protocol as Tan et al. (2023).
For the silica glass system, we also use the dataset from Tan et al. (2023), which consists of large SiO2 structures extracted from diverse, long molecular dynamics trajectories. These trajectories span a range of physical conditions, including temperature equilibration, cooling, uniaxial tension, compression, and more. The dataset is publicly available at: https://archive.materialscloud.org/records/tz6a0-hsb25. We adopt the same validation/test split protocol as Tan et al. (2023), and vary the size of the training set ($32$, $128$, and $1024$ samples).
Additional details regarding the DFT calculations used to generate both datasets are provided in the Supplementary Information of Tan et al. (2023).
@@
We follow the same experimental procedure as Tan et al. (2023): for prediction we compute the mean absolute error between predicted energy and forces $\hat{E},\bm{\hat{F}}$ against true ones:
@@
To evaluate uncertainty, we compute the Expected Calibration Error (ECE) on forces, following the approach proposed by Tan et al. (2023). This choice is motivated by the fact that forces are more sensitive to overfitting and distributional shift, and thus provide a more informative measure of epistemic uncertainty than energy-based metrics. We estimate uncertainties using the mean force per structure, denoted by $\bar{F}_{i}$ for structure $i$, which is obtained by averaging the per-atom forces and flattening across the spatial dimensions. The corresponding standard deviation, $\bar{\sigma}_{i}$, is computed as the square root of the average variance norm per atom, following the procedure in Tan et al. (2023). The ECE is then computed as:
@@
where $\mathbb{Q}$ is the gaussian quantile function. The integral is computed by the trapezoidal rule. We further evaluate the rank correlation between predicted variance $\bar{\sigma}_{i}^{2}$ and squared error norm per atom using the Spearman correlation coefficients using the Scipy (Virtanen et al., 2020) implementation.
B.3 MATPES Dataset
Dataset Availability:
We use the PBE version of the MATPES dataset (Kaplan et al., 2025), which is publicly available at https://matpes.ai/ . For the zero active learning (AL) step, the dataset is split as follows (fractions of the full dataset): initial training set =0.001=0.001, pool set =0.025=0.025, test set =0.79=0.79, and validation set =0.124=0.124. At each AL step, $10\%$ of candidate structures are selected from the pool loader. Selection is either random or based on the top $10\%$ most uncertain structures, where uncertainty is quantified as the standard deviation of the predicted energies.
Metrics:
We follow the same metrics as in the Silica experiment. Prediction accuracy is evaluated using the mean absolute error (MAE) for energies and forces:
where $i$ indexes the data points and $k$ indexes the spatial dimensions. For uncertainty quantification, we use the standard deviation of the predicted energies across stochastic samples.
$\sigma_{E}=\frac{1}{N-1}\sqrt{\sum_{i=1}^{N}\left(E_{i}-\hat{E}_{i}\right)^{2}}.$ (31)
@@ Appendix C Architectures and Training Details
All our models have been trained on a single 64GB A-100 Nvidia GPU. We report in Table 5 the hyperparameters used across experiments shared among different models. All models are trained on the training dataset and evaluated on the test dataset. The validation dataset is used to select the best checkpoint, corresponding to the smallest validation loss. The validation loss is computed using the same loss function as in training, but evaluated on the validation set.
Table 5: Summary of main training-related hyperparameters for different experiments. These hyperparameters are shared among different models. Adam is the optimiser in Kingma and Ba (2014), while AdamW the one in Loshchilov and Hutter (2017).
Hyper-parameter NBody Ammonia Silica MATPES
Optimization Epochs 10000 500 50 25 (5 AL steps)
Batch Size 100 64 8 128
Optimiser Adam Adam AdamW AdamW
Loss MSE MAE HuberLoss(0.01) HuberLoss(0.01)
Maximum learning rate 5e45e-4 1e31e-3 1e51e-5 1e51e-5
Learning rate scheduling Const. Reduce On Plateau Cosine Const.
Gradient Clip Norm Value NA NA 0.5 0.5
Warmup Iteration Phase NA NA 5% N
@@ C.1 NBody System
For the EGNN baseline, we employ the velocity variant from Satorras et al. (2021), where the initial velocity norm of each particle is mapped through a linear layer and included in the initial node features $\mathbf{h}^{0}_{i}$. Particle charges are again encoded as edge attributes via the pairwise product $a_{ij}=c_{i}c_{j}$. The message and update functions have the same parameterisation as in the GNN model. Following Satorras et al. (2021), no regression head is used, as the network directly predicts particle positions.
Regarding the baseline methods, for MC Dropout we follow Gal and Ghahramani (2016) and apply dropout with varying probabilities after all linear layers of the message-passing network. For Deep Ensemble, we train the same architecture four times with different random initialisation seeds.
Finally, for BLIP, we employ a compact invariant posterior network consisting of two linear layers of size $64$ with swish activations, used for both the message and update coefficients. Dropout probabilities $p$ are predicted and converted to coefficients via p/(1p)p/(1-p). The message inference network head takes as input the edge attribute $a_{ij}$ together with $\mathbf{h}^{0}_{i}$ and $\mathbf{h}^{0}_{j}$, whereas the update inference network head only takes $\mathbf{h}^{0}$. For stability, we clamp the maximum VAD coefficients to $4.0$. During training, we use a prior dropout probability of $0.5$ and weight the regularisation term $D_{\mathrm{KL}}$ by $\lambda=0.01$.
@@ C.2 Ammonia Molecule System
We adopt the PaiNN architecture (Schütt et al., 2021) as our machine learning interatomic potential. To ensure a fair comparison, we use the same hyperparameter settings as in Tan et al. (2023), which are publicly available at: https://github.com/learningmatter-mit/UQ_singleNN.
@@
The message inference network head takes as input the squared interatomic distance $\|\bm{r}_{i}-\bm{r}_{j}\|^{2}$, whereas the update inference network head takes only an embedding of the element atomic number $Z$. For stability, we clamp the VAD coefficients to $4.0$. During training, we use a prior dropout probability of $0.5$ and weight the regularisation term $D_{\mathrm{KL}}$ by $\lambda=10$.
@@ C.3 Silica Glass System
We adopt the ORB v3 architecture (Rhodes et al., 2025) as our machine learning interatomic potential, pretrained on ab initio molecular dynamics data from OMAT 24 (Barroso-Luque et al., 2024). The Deep Ensemble variant finetunes four different models with different initialisation seeds.
@@
The edge network takes as input a concatenation of the pairwise distance vector (normalised to unit length) and 20 Gaussian radial basis functions applied to the raw edge distance. The node network processes atomic number embeddings. For stability, we clamp the VAD coefficients to the range $10^{-5},0.11]$. During training, we use a prior dropout probability of $0.01$ and weight the regularisation term $D_{\mathrm{KL}}$ by $\lambda=0.001$. The lower clamp values, $\lambda$, and reduced prior dropout are chosen because we find that the original ORB model weights are highly sensitive to even minimal perturbations, likely because they were not trained with dropout or other weight-perturbation techniques. In addition, we use a different learning rate (1e41e-4) for the inference network to enable faster adaptation compared to the pretrained weights.
C.4 MATPES dataset
We employ the same hyperparameters as for the Silica Glass System.
@@ Appendix D Algorithms
12:  $\mathcal{L}\leftarrow\lambda_{\rm{E}}\text{MAE}_{\rm{E}}+\lambda_{\rm{F}}\text{MAE}_{\rm{F}}+\lambda D_{\mathrm{KL}}(\bm{\alpha},\bm{\beta})$ {Loss includes MAEs and KL divergence term from Eq. equation 10}
@@
Uncertainty quantification is performed via multiple stochastic forward passes through the model using Monte Carlo (MC) sampling. Importantly, the inference network is evaluated only once to extract the variational parameters $\bm{\alpha},\bm{\beta}$, which are reused across MC samples. This amortization significantly reduces computational overhead. While more accurate moment-matching approximations could be used for the output distributions, we leave this for future work. Finally, if only predictive means are needed and uncertainty estimates are not required, a single deterministic forward pass can be performed using the mean of the variational distributions (i.e., without stochastic sampling), effectively matching the performance speed of a standard MLIP.
@@
Appendix E Computational Cost Analysis
A central motivation of our approach is to obtain high-quality uncertainty estimates and mean predictions with minimal additional computational overhead. In this section, we quantify the training, inference, sampling and memory cost of our method and compare it against deep ensembles. In Table 6 we summarise relevant statistics as model parameters, training and inference FLOPs.
Training Cost:
Table 6: FLOPs breakdown for ORBv3-inf-omat24 (Base) model for training and inference.
[…]
Params 21.02 21.05 $21.02\times N$
Params increase from Base $\times 1$ $\times 1.001$ $\times N$
Training FLOPs / atom ($\times 10^{12}$) 1.65 3.1 $1.65\times N$
FLOPs increase from Base $\times 1$ $\times 1.9$ $\times N$
Inference FLOPs / atom ($\times 10^{12}$) 1.65 1.65 $1.65\times N$
FLOPs increase from Base $\times 1$ $\times 1$ $\times N$
Our method trains a single backbone model jointly with a lightweight inference network $E$ that contains only $p\%$ of the parameters of the main model (typically $0.1\%\text{ to }5\%$). Let $C_{\text{main}}$ denote the cost in FLOPs of training the base model, and $C_{\text{inf}}$ the one for the inference network. The total training cost of BLIP is therefore
$C_{\text{BLIP}}=C_{\text{main}}+C_{\text{inf}}\approx(2+\epsilon)\,C_{\text{main}},$
where $\epsilon=C_{\text{inf}}/C_{\text{main}}\approx p/100\ll 1$. The factor of two comes from the local reparameterization trick, which requires computing both the mean and variance. In contrast, an ensemble of $N$ independently trained models incurs a cost of
$C_{\text{ens}}=N\,C_{\text{main}},$
which is $N$-times larger than training a single model. For commonly used ensemble sizes ($N=5$), this results in roughly a $5\times$ increase in both computation and memory footprint.
Inference Cost:
At inference time, our method evaluates a single forward pass through the backbone as usual, since we set the variational dropout coefficient to zero. This approach does not add any overhead compared to the cost of evaluating the backbone model. On the other hand, an ensemble of $N$ models requires $N$ separate forward passes, thus scaling linearly in size.
Sampling Cost:
For performing a single sample from the posterior, the forward FLOPs required are the same as during training. Advanced techniques such as vectorial-mapping, moment matching or moment linearization could be employed to drastically reduce the computational cost of predictive distribution moment estimation, but we leave this for future work.